# Dust evolution during the protostellar collapse: influence on the coupling between the neutral gas and the magnetic field

### v1

@@ Appendix D Ambipolar drift intensity in a collapsing core
Figure 15: Face-on (left) and edge-on (right) slices of the ambipolar drift intensity $\delta_{\mathrm{AD}}$ extracted from a non-ideal MHD collapse model (mmMRNmhd, Lebreuilly et al. 2020) computed with RAMSES (Teyssier 2002) at $t=80\ \mathrm{Kyr}$.
@@ Appendix E Impact of grain-grain erosion and electrostatic repulsion on the Ohm resistivity
Figure 16: Ohm resistivity $\eta_{\mathrm{O}}$ evolution during the protostellar collapse for different values of the ambipolar drift intensity or equivalently different ambipolar drift velocities : $\delta_{\mathrm{AD}}=0.1\ (0.1V_{\mathrm{AD}})$, $\delta_{\mathrm{AD}}=1\ (V_{\mathrm{AD}})$, and $\delta_{\mathrm{AD}}=10\ (10V_{\mathrm{AD}})$. Grain-grain electrostatic repulsion (first column) and erosion (second column) are also included (dotted lines). Second row : likewise but with a sticking efficiency of 0.5.
@@
Similarly to the analysis carried out in Sect. 4.2.2, we investigate here the impact of grain-grain electrostatic repulsion and erosion on the Ohm resistivity $\eta_{\mathrm{0}}$ for different values of the ambipolar drift intensity $\delta_{\mathrm{AD}}$. Looking at the first column, we can see that contrary to the ambipolar resistivity, the Ohm resistivity remains almost unaffected at all densities even for a weak ambipolar drift intensity ($\delta_{\mathrm{AD}}=0.1$) for which electrostatic repulsion is expected to significantly inhibit small grains coagulation. While the dust contributes directly to the ambipolar resistivity $\eta_{\mathrm{AD}}$ at low density (any change in the small grain population reflects almost immediately on the resistivity profile), it only acts indirectly on the ohmic resistivity, via electron capture. At low densities, the cross-section for electron capture is very low and only a sufficient abundance of small grains, like the one permitted by a lower sticking efficiency ($\epsilon_{\mathrm{stick}}=0.5$, bottom left panel) entails an efficient collection of electrons and thus a rise in the ohmic resistivity. At high density, the grain collisional velocities are usually above the Coulomb velocity threshold (Eq. 18), thus suppressing the effect of electrostatic repulsion.
@@
When it comes to the impact of erosion (right column), it is clear that for low and intermediate values of ambipolar drift intensity ($\delta_{\mathrm{AD}}=0.1$ and $\delta_{\mathrm{AD}}=1$), the collisional velocities are not high enough to induce a sufficient replenishment of small grains prior to $n_{\mathrm{H}}=10^{12}\ \mathrm{cm^{-3}}$. However, beyond this density, the collisional velocities and dust size ratios $s_{\mathrm{ratio}}$ (see Eq. 17) are large enough to enhance the erosion efficiency and thus to allow the excess of small grains to efficiently capture the electrons and to allow a soaring of the Ohm resistivity. Nevertheless, in order to increase the resistivity at lower density, one needs to rely on a stronger ambipolar drift intensity $\delta_{\mathrm{AD}}=10$, as seen by looking at the purple dotted curve. Note that for $\delta_{\mathrm{AD}}=0.1$ and $\delta_{\mathrm{AD}}=1$, a sticking efficiency of $\epsilon_{\mathrm{stick}}=0.5$ does reduce the small grain coagulation rate, but is also detrimential to erosion, since only half of the collisions are destructive. This results in an overall suppression of the late rise in ohmic resistivity.
@@
In any case, the Ohm resistivity remains at least two orders of magnitude below the MRN reference because no process is capable of completely making up for the very short coagulation timescales. Besides, it remains dominated by the ambipolar resistivity $\eta_{\mathrm{AD}}$ at least up to the final density value of our simulations, i.e., $n_{\mathrm{H}}=2\times 10^{13}\ \mathrm{cm^{-3}}.$
@@
Appendix F $\gamma_{\mathrm{grain}}=190\ \mathrm{erg}\ \mathrm{cm^{-2}}$ : impact of the grain-grain sticking efficiency $\epsilon_{\mathrm{stick}}$
@@
Figure 17: Likewise Fig. 2 but investigating the grain sticking efficiency ($\gamma_{\mathrm{grain}}=190\ \mathrm{erg}\ \mathrm{cm^{-2}}$).
@@
We vary here the sticking efficiency of grains $\epsilon_{\mathrm{stick}}$ (included in our coagulation algorithm) and explore its impact on the evolution of the dust distribution as well as on the magnetic resistivities. Note that $\epsilon_{\mathrm{stick}}=0.5$ implies that half of the collisions result in bouncing rather than coagulation or fragmentation, i.e. that both grains remain intact. Fig. 17 indicates that a value of $0.1$ would significantly hinder the small grains coagulation at low density, allowing to recover the MRN resistivities to some extent. Note that a value of 0.1 means that only one collision out of ten leads to grain sticking, and thus to coagulation. A value of $0.3$ slows down the coagulation process (as readily seen on the ambipolar resistivity $\eta_{\mathrm{AD}}$ profile) and allows the ambipolar resistivity to be notably reduced prior to $n_{\mathrm{H}}=10^{9}\ \mathrm{cm^{-3}}$. However, for values beyond $\epsilon_{\mathrm{stick}}=0.5$, no significant changes are induced. This parameter also has a clear impact on the peak size of the dust distribution. Indeed, a lower sticking efficiency implies a longer coagulation timescale and consequently a reduced dust peak size.

### v2

@@ Appendix D Ambipolar drift intensity in a collapsing core
Figure 15: Face-on (left) and edge-on (right) slices of the ambipolar drift intensity, $\delta_{\mathrm{AD}}$, extracted from a non-ideal MHD collapse model (mmMRNmhd, Lebreuilly et al. 2020) computed with RAMSES (Teyssier 2002) at $t=80\ \mathrm{Kyr}$.
@@ Appendix E Impact of grain-grain erosion and electrostatic repulsion on the Ohm resistivity
Figure 16: Ohm resistivity, $\eta_{\mathrm{O}}$, evolution during the protostellar collapse for different values of the ambipolar drift intensity or equivalently different ambipolar drift velocities: $\delta_{\mathrm{AD}}=0.1\ (0.1V_{\mathrm{AD}})$, $\delta_{\mathrm{AD}}=1\ (V_{\mathrm{AD}})$, and $\delta_{\mathrm{AD}}=10\ (10V_{\mathrm{AD}})$. Grain-grain electrostatic repulsion (first column) and erosion (second column) are also included (dotted lines). Second row: Same, but with a sticking efficiency of 0.5.
@@
Similarly to the analysis carried out in Sect. 4.2.2, we investigated the impact of grain-grain electrostatic repulsion and erosion on the Ohm resistivity $\eta_{\mathrm{0}}$ for different values of the ambipolar drift intensity $\delta_{\mathrm{AD}}$. Looking at the first column, we can see that contrary to the ambipolar resistivity, the Ohm resistivity remains almost unaffected at all densities even for a weak ambipolar drift intensity ($\delta_{\mathrm{AD}}=0.1$) for which electrostatic repulsion is expected to significantly inhibit small grains coagulation. While the dust contributes directly to the ambipolar resistivity, $\eta_{\mathrm{AD}}$, at low density (any change in the small grain population is reflected almost immediately in the resistivity profile), it only acts indirectly on the ohmic resistivity, via electron capture. At low densities, the cross-section for electron capture is very low and only a sufficient abundance of small grains, such as the one enabled by a lower sticking efficiency ($\epsilon_{\mathrm{stick}}=0.5$, bottom left panel), allows for an efficient collection of electrons and, thus, a rise in the ohmic resistivity. At high density, the grain collisional velocities are usually above the Coulomb velocity threshold (Eq. 18), thereby suppressing the effect of electrostatic repulsion.
@@
When it comes to the impact of erosion (right column), it is clear that for low and intermediate values of ambipolar drift intensity ($\delta_{\mathrm{AD}}=0.1$ and $\delta_{\mathrm{AD}}=1$), the collisional velocities are not high enough to induce a sufficient replenishment of small grains prior to $n_{\mathrm{H}}=10^{12}\ \mathrm{cm^{-3}}$. However, beyond this density, the collisional velocities and dust size ratios $s_{\mathrm{ratio}}$ (see Eq. 17) are large enough to enhance the erosion efficiency and thus to allow for the excess of small grains to efficiently capture the electrons and to allow for a soaring in terms of the Ohm resistivity. Nevertheless, in order to increase the resistivity at lower density, we need to rely on a stronger ambipolar drift intensity $\delta_{\mathrm{AD}}=10$, as seen by looking at the purple dotted curve. We note that for $\delta_{\mathrm{AD}}=0.1$ and $\delta_{\mathrm{AD}}=1$, a sticking efficiency of $\epsilon_{\mathrm{stick}}=0.5$ does reduce the small grain coagulation rate, but is also detrimential to erosion, since only half of the collisions are destructive. This results in an overall suppression of the late rise in ohmic resistivity.
@@
In any case, the Ohm resistivity remains at least two orders of magnitude below the MRN reference because no process is capable of completely making up for the very short coagulation timescales. Besides, it remains dominated by the ambipolar resistivity $\eta_{\mathrm{AD}}$ at least up to the final density value of our simulations, namely, $n_{\mathrm{H}}=2\times 10^{13}\ \mathrm{cm^{-3}}.$
@@
Appendix F Impact of the grain-grain sticking efficiency, $\epsilon_{\mathrm{stick}}$
@@
Figure 17: Same as Fig. 2, but investigating the grain sticking efficiency ($\gamma_{\mathrm{grain}}=190\ \mathrm{erg}\ \mathrm{cm^{-2}}$).
@@
Here, we vary the sticking efficiency of grains, $\epsilon_{\mathrm{stick}}$, (included in our coagulation algorithm) and explore its impact on the evolution of the dust distribution as well as on the magnetic resistivities. We note that $\epsilon_{\mathrm{stick}}=0.5$ implies that half of the collisions result in bouncing rather than coagulation or fragmentation, namely, that both grains remain intact. Figure 17 indicates that a value of $0.1,$ which would significantly hinder the small grains coagulation at low density, allowing us to recover the MRN resistivities to some extent. We note that a value of 0.1 means that only one collision out of ten leads to grain sticking, thus leading to coagulation. A value of $0.3$ slows down the coagulation process (as readily seen on the ambipolar resistivity $\eta_{\mathrm{AD}}$ profile) and allows for the ambipolar resistivity to be notably reduced prior to $n_{\mathrm{H}}=10^{9}\ \mathrm{cm^{-3}}$. However, for values beyond $\epsilon_{\mathrm{stick}}=0.5$, no significant changes are induced. This parameter also has a clear impact on the peak size of the dust distribution. Indeed, a lower sticking efficiency implies a longer coagulation timescale and, consequently, a reduced dust peak size.